OpenStudy (anonymous):

A sphere of radius 9 is centered at the point (-3,-2,-1). What are the coordinates of the two points where the surface of the sphere intersects the Z-axis?

OpenStudy (anonymous):

sphere equation : (x+3)^2 + (y+2)^2 + (z+1)^2 = 81 z intersection means both Y and X equal zero plug it into the equation gives : 9 + 4 + z^2 + 2z + 1 = 81 z^2 + 2z - 67 = 0 quadratic formula gives you : 7.246 -9.246